February 1999: Installation of the Modecleaner Tank TCMa
March 1999: Inside Modecleaner Tank TCMa
March 1999: Inside Modecleaner Tank TCMb
Impressions from the Site in Spring
May 1999: Breadboard and Scanner
June 1999: Test Run of the Main Suspension
Meanwhile at the Site
December 1999: Status Report
Both mode-cleaners (MC) are installed and lock reliable to the light of a prestabilized NPRO (700 mW). Only little time was spent on the mode matching, hence we got only 85% of the light into the first MC. The second MC sees a nicely spatial filtered beam, and the reflected power goes down to 5% once it is locked. The lock bandwidth of MC1 is above 1 kHz; at the moment MC2 is locked with only 400 Hz unity gain.
The first main triple pendulum suspension system with an Al dummy mirror is installed in TFE. This pendulum is damped by means of a 6-channel local control board with an attached digital board. So it is possible to read the flag position and the feedback signal with a LabView program which can be used to change gains and offsets as well. Mike and the suspension team are in the process of installing the first double-triple pendulum for MCE. It will be pre-assembled outside of TCE and installed early next year.
Power Recycling Tank
All parts are machined and cleaned. The pendulums for MPR and BDIPR will be installed once the TCE pendulums are ready.
The central cluster (without TCOa and TCOb) is assembled and attached to the mode-cleaner section. The pressure in the central cluster is 3x10-8 mbar. As the concentration of hydrocarbons is a little high in the mode-cleaner section an aperture was installed at the output of TCMb in order to form a differential pumping system for hydrocarbons between the mode-cleaner section and the main cluster. After the installation of the MFE pendulum we attached the east end tank TFE to the gate valve and pumped it down. The residual gas spectrum looks quite promising.
The first version of the "LabView Control Program Backbone" (for experts: the MCPs, data socket server, and handys) was installed and works fine.
Not much happened concerning the DAQ System in Ruthe, but a lot of preparation work has been done. We expect a lot of progress next year. The new radio link (34 Mb/s) will be installed until the end of February. The first data chain check from the site to Hannover and an overnight transfer to Potsdam is planned for March.